# Likelihood and appearance of life beyond the Earth: An astronomical perspective

### v1

@@ Acknowledgements.
Floris van der Tak is a Senior Scientist at SRON Netherlands Institute for Space Research, and Honorary Professor of Astrochemistry and Exoplanet Habitability at the University of Groningen. He is a Board member of the Netherlands Origins Center, and Scientific Coordinator of the EU-funded interdisciplinary research program oLife to investigate the origin of life on Earth and other planets. As of 2023, he has published 240 articles in peer-reviewed journals, and supervised 54 students and postdocs. Before his current position, he studied and worked in Bonn (Germany), Leiden (The Netherlands), and Berkeley (United States). His research focuses on the formation of stars from interstellar gas clouds, and the habitability of exoplanets. His students and postdocs use radio and infrared telescopes on Hawaii, in Chile, and in space.

### v2

@@ Acknowledgements.
Floris van der Tak is a Senior Scientist at Space Research Organisation Netherlands (SRON), and Honorary Professor of Astrochemistry and Exoplanet Habitability at the University of Groningen. He is a Board member of the Netherlands Origins Center and the PRELIFE consortium, and Scientific Coordinator of the EU-funded interdisciplinary research programs oLife and Evolve to investigate the origin of life on Earth and other planets. As of 2025, he has published 250 articles in peer-reviewed journals, and supervised 77 students and postdocs. Before his current position, he studied and worked in Bonn (Germany), Leiden (The Netherlands), and Berkeley (United States). His research focuses on the formation of stars from interstellar gas clouds, and the habitability of exoplanets. His students and postdocs use radio and infrared telescopes on Hawaii, in Chile, and in space.